### v2

# Efficient Indirect LLM Jailbreak via Multimodal-LLM Jailbreak

Abstract
This paper focuses on jailbreaking attacks against large language models (LLMs), eliciting them to generate objectionable content in response to harmful user queries. Unlike previous LLM-jailbreak methods that directly orient to LLMs, our approach begins by constructing a multimodal large language model (MLLM) built upon the target LLM. Subsequently, we perform an efficient MLLM jailbreak and obtain a jailbreaking embedding. Finally, we convert the embedding into a textual jailbreaking suffix to carry out the jailbreak of target LLM. Compared to the direct LLM-jailbreak methods, our indirect jailbreaking approach is more efficient, as MLLMs are more vulnerable to jailbreak than pure LLMs. Additionally, to improve the attack success rate of jailbreak, we propose an image-text semantic matching scheme to identify a suitable initial input. Extensive experiments demonstrate that our approach surpasses current state-of-the-art jailbreak methods in terms of both efficiency and effectiveness. Moreover, our approach exhibits superior cross-class generalization abilities. The code is available here.
1 Introduction
Recently, large language models (LLMs) such as ChatGPT [7] have been widely deployed. These models exhibit advanced general abilities but also pose serious safety risks such as truthfulness, toxicity, and bias [17, 26, 34, 3, 9]. Typically, there exists a type of attack called jailbreaking attack, which can elicit LLMs to generate objectionable content in response to users’ harmful queries. For example, a pioneering work [45] has found that a specific prompt suffix allows the jailbreak of most popular LLMs. However, the efficiency of those methods is recognized to be relatively low, primarily attributed to the challenges of discrete optimization in finding the textual jailbreaking suffix (JBtxt).
On the other hand, there is a surge of interest in multimodal large language models (MLLMs), which enable users to provide images as input to LLMs. [21, 14, 44, 11, 4, 43, 6, 37, 1]. Consequently, research on jailbreak has been extended from LLMs to MLLMs. Furthermore, it has been demonstrated that performing MLLM-jailbreak is easier and more efficient than performing LLM-jailbreak [31, 28]. This is largely because finding jailbreaking images within continuous pixel spaces is significantly easier and more flexible than identifying jailbreaking text within discrete token spaces.
Inspired by that, we propose an efficient indirect LLM-jailbreaking approach that constructs an MLLM built upon the target LLM and subsequently performs MLLM-jailbreak. Specifically, we adopt a widely used LLM-jailbreaking strategy [45], which seeks to identify a specific prompt suffix, denoted as JBtxt. When appended to harmful queries, this JBtxt suffix can elicit LLMs to generate objectionable content. Our contribution lies in efficiently finding such JBtxt by leveraging the MLLM-jailbreak process. As shown in Fig.1, the workflow of our approach is as follows: (1) Given a target LLM to be jailbroke (e.g., LLaMA2), we first construct an MLLM by integrating a visual module into the target LLM. (2) We then perform an efficient MLLM-jailbreak. (3) Instead of using the jailbreaking image (referred to as JBimg), we obtain the output features from the visual module—namely, the Jailbreaking embeddings (JBemb)—and convert them into textual strings using our De-embedding and De-tokenization operations. (4) We regard these textual strings as JBtxt and append them to the harmful queries to carry out the jailbreak of the target LLM.
Essentially, our approach leverages an MLLM-jailbreak to achieve the ultimate goal of LLM jailbreak. This indirect jailbreaking scheme offers flexibility for both white-box and black-box jailbreak. In the context of white-box jailbreak, the process of converting JBemb to JBtxt allows us to generate a set of candidate JBtxt. Unlike GCG [45], which produces only a single JBtxt, our approach outputs a diverse set of high-quality JBtxt, significantly enhancing the attack success rate (ASR) of jailbreak.
Black-box jailbreak is preferred in real-world applications. In our approach, the transition from MLLM-jailbreak to LLM-jailbreak enables a flexible and effective black-box jailbreak strategy. Specifically, we can construct the MLLM with a surrogate LLM and obtain the JBtxt through MLLM-jailbreak. Subsequently, the JBtxt can be transferred to jailbreak the target LLM in a black-box manner.
Regarding the MLLM-jailbreak, we observe that the ASR of jailbreak is closely related to the initial input image (namely JBinit). By selecting an appropriate JBinit, we can significantly enhance the MLLM-jailbreaking ASR. To achieve this, we propose an image-text semantic matching scheme to identify this suitable JBinit. Our scheme attempts to align the embedding of JBinit with that of harmful queries. The underlying intuition is that aligned embeddings between the image and text inputs are more effective at triggering the model’s cross-attention mechanism, thereby exerting a stronger influence on the answer generation process. As a result, our JBinit can effectively shift the model’s response from “Sorry, I cannot" to “Sure, here is", thus achieving a successful jailbreak.
Our approach is also related to embedding-based LLM jailbreak methods [42], which focus on direct optimizing token embeddings. Compared to token-based methods such as GCG, these approaches are more efficient, as they rely on continuous optimization. However, previous embedding-based jailbreaks have been shown to be largely ineffective, primarily because the obtained embeddings often lack corresponding tokens [45]. In contrast, our method does not directly optimize JBemb. Instead, we indirectly obtain JBemb by optimizing JBimg. This procedure can be viewed as regularizing the optimization of JBemb through the use of the visual module. Since the visual module (e.g., a CLIP encoder) is trained with an image-text alignment objective, it increases the likelihood that our JBemb corresponds to valid tokens. Consequently, our approach surpasses traditional embedding-based jailbreak methods in terms of effectiveness.
Regarding the evaluation of LLM jailbreak, previous methods typically rely on benchmark datasets that mix various types of harmful behaviors. In contrast, we propose to categorize them into fine-grained classes, such as violence, financial crimes, cyber crimes, drug crimes, etc. This enables us to assess the cross-class generalization ability of jailbreaks—specifically, whether JBtxt generated from one class can effectively transfer to jailbreak other classes.
We conduct extensive experiments demonstrating that our approach outperforms current state-of-the-art jailbreaking methods in both efficiency and effectiveness. Efficiency is a key advantage of our approach. For instance, for the same jailbreaking tasks, our method requires only approximately 3% of GCG’s runtime. Furthermore, our approach provides notable flexibility for black-box jailbreak, enabling successful jailbreaks of recent LLMs such as Mistral-v0.3, Gemma2, Deepseek-R1, Grok-3, and DouBao1.5. Furthermore, it exhibits superior cross-class generalization ability of jailbreak. This finding suggests that improving the ASR for a specific class can benefit not only from data within the class itself, but also from data drawn from semantically related classes.
2 Related Work
LLM-jailbreak. Prior work on LLM jailbreak primarily focuses on crafting input prompts—referred to as jailbreaking prompts—to circumvent model alignment constraints. Early approaches relied on manually designing prompts with specific strategies, such as role-playing [26], prompt injection, and privilege escalation [33]. More recently, research has shifted toward automating the generation of jailbreak prompts to reduce the reliance on manual efforts.
Existing jailbreak techniques can be broadly categorized into two types. The first is token-based jailbreak, which directly optimizes discrete token sequences. For instance, [16] leverage time-based SQL injection techniques to automatically generate jailbreak prompts. [22] introduce autoDAN, a hierarchical genetic algorithm that evolves stealthy and effective prompts. The GCG method proposed by [45] appends specific suffixes to input queries to enable automated jailbreaks.
The second category is embedding-based jailbreak, which starts by optimizing token embeddings and subsequently projects them back into discrete tokens. A notable challenge with this approach is that the optimized embeddings often do not correspond to any valid token. To address this, Prompts Made Easy (PEZ) [42] adopts a quantized optimization strategy that adjusts continuous embeddings while ensuring they can be effectively mapped back to tokens.
MLLM-jailbreak. Recently, the safety risks posed by multimodal large language models (MLLMs) have attracted growing concern. Several jailbreak techniques specifically designed for MLLMs have emerged. For instance, a pioneering study by Carlini [10] demonstrated that adversarial prompts in the form of continuous-domain images can effectively induce MLLMs to generate harmful or toxic content. Similarly, [32] proposed decomposing a malicious prompt into a benign textual query combined with malicious triggers—either OCR-based textual triggers or visual triggers—exploiting the joint embedding space to enable the jailbreak. These methods collectively reveal that MLLM jailbreaking achieves significantly higher success rates, often surpassing those of traditional LLM jailbreak strategies.
3 Our Approach
Unlike previous LLM-jailbreak methods that directly orient to LLMs, our approach first constructs a multimodal large language model (MLLM) by incorporating a visual module. We then leverage an efficient MLLM-jailbreak to ultimately achieve the goal of LLM-jailbreak. In the following sections, we first provide an overview of our framework, followed by a detailed description of each component.
3.1 Overview
Our approach consists of four steps, as illustrated in Fig. 2. In Step 1, we construct an MLLM by incorporating a visual module and perform an MLLM-jailbreak to obtain the jailbreaking embedding, denoted as JBemb. In Step 2, the obtained JBemb is converted into textual form, referred to as JBtxt. Finally, in Step 3, we use these JBtxt to carry out the LLM-jailbreak.
Our empirical study reveals that the Attack Success Rate (ASR) of jailbreak is sensitive to the initial input image (JBinit) used to initialize the MLLM-jailbreaking process. To address this, we introduce an image-text semantic matching scheme as Step 0 to identify a suitable JBinit.
Figure 1: The indirect jailbreaking scheme of our approach begins with constructing an MLLM by integrating a visual module. We then perform an efficient MLLM-jailbreak to obtain the JBemb, which is subsequently converted into JBtxt for jailbreaking the target LLM.
3.2 MLLM Construction
Given a target LLM, we aim to construct a multimodal LLM (MLLM) that incorporates the target LLM as its core. This can be achieved through two approaches: building and training the MLLM ourselves or utilizing an existing model.
For the first approach, we adopt a standard pipeline that involves integrating a visual module into the target LLM. We then fine-tune the MLLM using a suitable multimodal dataset. Crucially, during the fine-tuning process, the LLM itself is frozen, and only the visual module is updated. This design ensures that the LLM embedded within the MLLM remains identical to the target LLM, thereby enabling effective transferability from MLLM-jailbreak to LLM-jailbreak.
In practice, however, it is often expensive to construct and train an MLLM from scratch. Thanks to the HuggingFace open-source platform, most popular LLMs already have corresponding MLLMs couple. For each LLM-MLLM pair, all intermediate models are systematically organized into a Model-Tree. Given a target LLM, one can simply identify the corresponding MLLM from the Model-Tree and proceed directly to the MLLM-jailbreaking stage.
3.3 MLLM Jailbreak
We adopt a strategy for MLLM-jailbreak analogous to that used in LLM-jailbreak, i.e., encouraging the MLLM to generate a response that begins with a positive affirmation such as “Sure, here is a (content of query)." Our method is inspired by adversarial attacks, which typically perturb visual inputs to manipulate a model’s output. However, jailbreak fundamentally differs from adversarial attacks in that it targets a generative task rather than a discriminative one. To bridge this gap, we adapt adversarial techniques by replacing the conventional cross-entropy loss with a maximum likelihood objective, thereby encouraging the MLLM to produce affirmative responses. This objective directly maximizes the likelihood of the desired output, making it better suited for the generative nature of jailbreak.
Specifically, for each harmful query $q_{i}$, we provide a corresponding target answer $a_{i}$, thereby creating a dataset of harmful behaviors $B=\{(q_{i},a_{i}),i=0,...,N\}$. We then formulate the MLLM-jailbreak as the process of finding a perturbation $\delta$ that encourages the generation of the target answer $a_{i}$ when the model is presented with the harmful query $q_{i}$, as follows,
$\mathop{\max}_{\delta}\sum_{i=0}^{M}log(p(a_{i}|q_{i},\widetilde{x}))$ (1)
[…]
$\qquad||\delta||_{p}<\epsilon$
where $p(a_{i}|q_{i},\widetilde{x})$ represents the likelihood of the MLLM generating the answer $a_{i}$ when provided with the image $\widetilde{x}$ and the text query $q_{i}$. Here, $\epsilon$ denotes the attack budget, and $\widetilde{x}$ refers to the jailbreaking image (JBimg). The optimization process is performed over a set of $M$ query-answer pairs $\{(q_{i},a_{i}),i=0,...,M\}$. This problem can be effectively solved by adapting the Projected Gradient Descent (PGD) algorithm [24].
3.4 LLM Jailbreak
After completing MLLM-jailbreak, we proceed to the next stage: carry out LLM-jailbreak. At this stage, the key is to derive a textual prompt suffix JBtxt, rather than directly utilizing the output of MLLM-jailbreak, i.e., JBimg.
Note that it is the output of the visual module—serving as the input to the LLM—that facilitates the success of jailbreak. We refer to this output as JBemb, a sequence of continuous embeddings that facilitates LLM-jailbreak. Consequently, transitioning from MLLM-jailbreak to LLM-jailbreak entails converting JBemb back into the text space, resulting in the textual JBtxt.
In our approach, we propose De-embedding and De-tokenization operations to convert JBemb into JBtxt. These operations are intended to reverse the Embedding and Tokenization procedures used in LLMs. Specifically, the embedding operation in LLMs maps each discrete token $t$ to its corresponding embedding vector $e$ via a token-embedding dictionary $(t,e)$. Accordingly, our De-embedding operation aims to invert this process by mapping a continuous embedding vector back to its most likely discrete token. This is achieved through a nearest-neighbor search across the embedding dictionary.
For each embedding vector $e_{l}$ in the sequence $(e_{1},...,e_{L})$, we identify the top-$K$ closest embedding vectors $\hat{e}^{k}_{l}$, where $k=1,...,K$. Performing this for all $e_{l}$ results in a $K\times L$ embedding pool $\{\hat{e}_{l}^{k}\}_{k=1,l=1}^{K,L}$ and a corresponding $K\times L$ token pool $\{\hat{t}_{l}^{k}\}_{k=1,l=1}^{K,L}$. The De-tokenization operation then converts these tokens into natural language words, yielding a $K\times L$ word pool $\{\hat{w}_{l}^{k}\}_{k=1,l=1}^{K,L}$. Finally, we randomly sample several word sequences from this word pool—each sampled sequence constituting a candidate JBtxt.
It is worth noting that we select the top-K nearest embeddings rather than solely the top-1. This strategy enables us to generate multiple JBtxt candidates instead of a single one. Moreover, we observe that each JBtxt has a certain probability of successfully achieving jailbreak. By ensembling these high-quality JBtxt candidates, we can substantially enhance the overall jailbreaking performance.
When performing black-box jailbreak, we first construct an MLLM based on a surrogate LLM model (e.g., LLaMA2), generate the corresponding JBtxt, and then transfer this JBtxt to effectively jailbreak the target LLM (e.g., Deepseek-R1).
Figure 2: The full workflow of our approach. Before Step1, we propose an image-text matching scheme to identify an appropriate initial input JBinit at Step0.
3.5 Finding an Appropriate JBInit
We observe that initializing the MLLM-jailbreak algorithm with a suitable starting point, namely JBInit, can substantially enhance the ASR of the MLLM-jailbreak. To this end, we propose an image-text semantic matching scheme. Specifically, our goal is to ensure that the embedding of JBInit is close to the embedding of the harmful queries. The rationale behind this scheme is that, for MLLM-jailbreak to succeed, the JBInit must exert a substantial influence on the LLM’s answer generation process. Particularly, it needs to turn a typical refusal response like “Sorry, I cannot" into a positive affirmation such as “Sure, here is". By ensuring that the embedding of JBInit closely aligns with that of the query, we can effectively trigger the model’s cross-attention mechanism, thereby exerting a stronger influence on the answer generation process.
Specifically, our image-text semantic matching scheme employs a network architecture akin to the CLIP model, as illustrated in Fig.2. Since the CLIP visual encoder serves as the visual module within the MLLM, our matching network reuses this same visual encoder. Additionally, we incorporate the CLIP text encoder to process textual inputs.
For the text input, we randomly select several keywords closely associated with a particular harmful class. For example, in the case of Weapons Crimes, these keywords might include firearms, illegal weapons, guns, etc. We then construct descriptive phrases from these keywords, such as “a photo of firearms,” which are better suited for encoding by the CLIP text encoder.
Regarding the image input, we initially employ an image search engine to retrieve candidate images using the selected keywords as search queries. These images are then ranked based on their CLIP similarity scores relative to the harmful phrases. The highest-ranked image is chosen as the input for the visual encoder.
Following this, we introduce an optimization algorithm designed to further modify the input image, with the objective of minimizing the distance between its embedding and that of the harmful phrases. Let’s denote the harmful phrases as $\{t_{i}\}_{i=0}^{N}$ and the highest-ranked image as $x$. Our matching algorithm seeks to find a perturbation $\Delta$ to be added to the image $x$ in order to maximize the image-text similarity score. Formally, this can be expressed as:
$\mathop{\min}_{\Delta}\sum_{i=0}^{N}log(\mathcal{L}_{\text{CLIP}}(\text{Enc}_{V}(\widetilde{x}),\text{Enc}_{L}(t_{i}))$ (2)
[…]
where $\text{Enc}_{V}$ is the CLIP visual encoder, $\text{Enc}_{L}$ is the CLIP text encoder, and $\widetilde{x}$ is the JBInit image obtained after perturbation.
By solving Eq( 2), we obtain an appropriate JBInit, which serves as an effective initialization point for MLLM-jailbreak.
4 Evaluation
Datasets. Several datasets have been introduced for evaluating jailbreak attacks, such as AdvBench [45], which includes a diverse range of harmful behaviors, including violence, financial crimes, drug-related crimes, and more. However, existing evaluation protocols often aggregate all these behaviors into a single pool during assessment.
In contrast, we advocate for a more fine-grained evaluation strategy by categorizing harmful behaviors into distinct classes and assessing jailbreak success for each class individually. This approach reveals that certain classes are inherently more resistant to jailbreak attempts than others.
Moreover, we can also assess the cross-class transferability of jailbreaks, which involves evaluating whether the JBtxt generated for one class can effectively jailbreak LLMs across other classes.
Specifically, we categorize AdvBench dataset into nine classes: “unlawful violence", “financial crimes", “property crimes", “drug crimes", “weapons crimes", “cyber crimes", “hate", “Suicide and Self-Harm", and “Fake info". Furthermore, since our approach involves MLLM-jailbreak, we extend the AdvBench dataset into a multimodal version, termed AdvBench-M, by augmenting it with images. Specifically, for each harmful class, we retrieve 20 semantically relevant images from the Internet using the Google search engine. We then employ CLIP ViT-L/14 [30] to select the top 10 images that best align with the queries associated with each harmful class.
In addition to AdvBench, we conducted extensive experiments on HarmBench dataset [25]. HarmBench consists of 400 harmful behavior instances. Notably, HarmBench encompasses several distinct behavior categories not included in AdvBench, such as copyright violations. To ensure a rigorous evaluation of our proposed method, we strictly followed the evaluation protocols outlined in the HarmBench paper.
Test Models. We evaluate our approach with both white-box and black-box jailbreaking settings. For the white-box setting, we conduct evaluations on open-source models, including LLaMA2-Chat-7B [40], InternLM2.5-7B [8], and Qwen2.5-3B-Instruct [39]. Specifically, for LLaMA2-Chat-7B, we utilize MiniGPT-4 [44] as the corresponding MLLM; for InternLM2.5-7B, we utilize InternVL2.0-8B [13] as the corresponding MLLM; and for Qwen2.5-3B-Instruct, we utilize InternVL2.5-4B [12] as its corresponding MLLM.
For the black-box setting, we generate the JBtxt by using InternVL2.0 as MLLM, and subsequently employ these textual suffixes to jailbreak closed-source models, including Mistral-v0.3 [19], Gemma2 [36], ChatGLM4 [18], Deepseek-R1 [15], Grok-3 [2], and DouBao1.5 [35].
Assessment. Evaluating the success of jailbreak is inherently challenging. The objective of jailbreak is to elicit the LLM to produce any response that complies with a harmful query, making it difficult to define a specific ground truth for each query. Moreover, we have observed cases where responses begin with seemingly affirmative phrases (e.g., “Sure, here is a…") but ultimately decline to fulfill the query. Consequently, prior methods often rely on manual inspection to judge jailbreak success. To overcome this limitation, we propose using the LLaMA Guard 2 tool [38] to automatically assess the success of jailbreak attempts.
Table 1: White-box jailbreak of LLaMA2 (via MiniGPT4-Jailbreak) on AdvBench-M dataset. Performance is evaluated using ASR (%).
Class AutoP AutoDAN SoftP GCG AdvA Ours
$D_{tr}$ $D_{t}$ $D_{tr}$ $D_{t}$ $D_{tr}$ $D_{t}$ $D^{o}_{t}$ $D_{tr}$ $D_{t}$ $D^{o}_{t}$ $D_{tr}$ $D_{t}$ $D^{o}_{t}$ $D_{tr}$ $D_{t}$ $D^{o}_{t}$
Unlawful violence 66 50 23 18 20 44 36 80 83 78 53 60 44 100 94 85
Financial crimes 80 93 33 46 40 32 37 100 89 81 60 0 50 93 98 74
Property crimes 73 84 17 48 80 62 36 100 98 97 60 68 55 93 98 69
Drug crimes 73 80 33 23 53 40 41 100 80 86 60 53 44 80 100 77
Weapons crimes 40 66 17 9 33 33 40 73 66 66 73 55 38 80 88 78
Cyber crimes 73 86 73 57 53 53 32 100 66 87 44 58 66 93 94 72
Hate 73 66 43 20 20 6.6 42 53 66 80 13 53 38 73 60 81
Suicide/Self-Harm 73 16 20 30 33 22 42 53 38 80 20 46 50 60 44 78
Fake info 73 93 90 40 33 6.6 30 80 86 66 93 33 44 93 86 93
Average ASR 69.3 70.4 38.8 32.3 40.6 33.2 37.3 82.1 74.7 80.1 52.9 47.3 47.7 85.0 84.7 78.6
Running Time 3.3h 2.1h 0.32h 11.2h 0.78h 0.37h
Table 2: White-box jailbreak of InternLM2.5 (via InternVL2.0-Jailbreak) on AdvBench-M dataset.
Class AutoP AutoDAN SoftP GCG AdvA Ours
$D_{tr}$ $D_{t}$ $D_{tr}$ $D_{t}$ $D_{tr}$ $D_{t}$ $D^{o}_{t}$ $D_{tr}$ $D_{t}$ $D^{o}_{t}$ $D_{tr}$ $D_{t}$ $D^{o}_{t}$ $D_{tr}$ $D_{t}$ $D^{o}_{t}$
Unlawful violence 6 20 20 11 46 38 60 60 53 47 40 44 64 94 80 86
Financial crimes 13 6 46 53 80 69 40 60 78 41 93 87 53 100 87 80
Property crimes 20 40 40 51 80 71 46 80 81 53 93 89 64 66 73 80
Drug crimes 26 33 20 60 60 53 53 67 37 35 67 67 53 86 93 100
Weapons crimes 33 13 20 38 53 61 53 73 68 47 46 55 64 88 100 66
Cyber crimes 40 26 62 60 86 80 60 53 35 37 73 90 53 86 96 66
Hate 13 20 26 40 13 0 60 46 20 41 46 13 70 86 66 73
Suicide/Self-Harm 26 6 26 11 40 33 66 20 22 29 67 44 70 66 66 86
Fake info 33 40 13 33 40 73 66 33 53 52 26 33 64 73 100 80
Average ASR 23.3 22.7 30.3 39.7 55.3 53.1 54.9 54.7 49.7 42.4 61.2 58.0 61.7 82.8 84.6 79.7
Running Time 4.2h 1.8h 0.40h 6.5h 1.05h 0.42h
4.1 White-box Jailbreak
We compare our approach with both token-based jailbreaking methods (i.e., AutoPrompt[27], AutoDAN [23], GCG [45], AdvAttack [5]) and embedding-based methods (i.e., PEZ [42], Soft Prompting [29]). Generally, the first category exhibits strong performance. GCG directly optimizes the JBtxt, while AutoDAN extends GCG by enhancing the readability of the generated JBtxt. AdvAttack introduces an adaptive strategy combining “Prompt + Random Search (RS) + Self-transfer," where the self-transfer component is employed to further boost jailbreaking effectiveness.
We conduct per-class evaluations on the AdvBench-M dataset. Specifically, for each class, we randomly sample 15 queries as the training set $D_{train}$ for learning the JBtxt, while the remaining queries constitute the testing set $D_{test}$.
Furthermore, to evaluate cross-class jailbreaking capability, we define $D_{test}^{other}$ as the union of all testing queries from classes other than the training class. For instance, we learn a JBtxt on Class 1’s $D_{train}^{1}$, and evaluate its performance both on the in-class test set $D_{test}^{1}$, and on the out-of-class aggregate set $D_{test}^{other}=D_{test}^{2}\cup D_{test}^{3}\dots\cup D_{test}^{9}$.
4.1.1 Compare to Token-based Jailbreak
We adopt Attack Success Rate (ASR) as the primary evaluation metric, computing it separately for each class on its corresponding $D_{train}$, $D_{test}$, and $D_{test}^{other}$. Tables 1 and 2 present the jailbreaking performance of LLaMA2 and InternLM2.5 on the AdvBench-M dataset. Due to space limitations, the results for jailbreaking Qwen2.5 are provided in the supplementary material.
From Tables 1 and 2, we observe that our approach consistently outperforms all baselines across $D_{train}$, $D_{test}$ and $D_{test}^{other}$. On LLaMA2, our method demonstrates a substantial improvement over the second-best performer, GCG, with an average ASR increase of 2.9 percentage points on $D_{train}$ (85.0% vs. 82.1%) and a more significant 10.0 percentage point improvement on $D_{test}$ (84.7% vs. 74.7%). When evaluating on InternLM2.5, the performance advantage of our approach becomes even more pronounced, outperforming the second-best competitor, AdvAttack, by 21.6 percentage points on $D_{train}$ (82.8% vs. 61.2%) and by 26.6 percentage points on $D_{test}$ (84.6% vs. 58.0%).
Notably, classes such as “Suicide/Self-Harm” and “Hate” remain particularly challenging to jailbreak. For instance, on LLaMA2, our method achieves only 44% and 60% ASR on $D_{test}$ for these classes respectively, and GCG reaches 38% and 66%. We attribute this difficulty to the fact that current models have undergone more intensive alignment efforts specifically for these sensitive classes. In contrast, certain classes demonstrate high vulnerability across most jailbreaking methods. For example, “Financial Crimes,” “Property Crimes,” and “Cyber Crimes” consistently yield high ASR across methods, with our approach achieving 98%, 98%, and 94% respectively on $D_{test}$ for LLaMA2. This disparity reveals the inherent imbalance in current safety alignment mechanisms across different harm classes. These findings not only expose vulnerable areas in existing safety guardrails but also provide concrete guidance for more comprehensive and robust safety alignment in future LLM development.
Efficiency. Our approach demonstrates markedly superior computational efficiency compared to token-based methods. Specifically, we compared the running time with GCG and found that GCG requires 11.2 hours to generate a single JBtxt for one harmful class, whereas our method completes the same task in just 0.37 hours—approximately 3% of GCG’s runtime. This represents a substantial reduction in computational overhead.
Table 3: White-box Jailbreak of LLaMA2 (via MiniGPT4-Jailbreak) and InternLM2.5 (via InternVL2.0-Jailbreak) on HarmBench dataset.
Model Dataset Attack Success Rate (%)
AutoP AutoDAN SoftP GCG AdvA Ours
LLaMA2 ALL 0 0.5 16.0 32.0 67.5 72.0
TEST 0 0.0 16.5 31.9 66.2 71.8
VAL 0 2.5 13.7 35.0 72.5 72.5
InternLM2.5 ALL 0 20.2 14.0 29.2 61.2 63.0
TEST 0 18.7 13.4 27.5 59.3 64.0
VAL 0 26.2 16.2 36.2 68.7 59.0
4.1.2 Compare to Embedding-based Jailbreak
Compared to token-based jailbreak, embedding-based jailbreaking methods can leverage straightforward and efficient continuous optimization, since token embeddings are continuous variables. This technique is often referred to as “soft prompting"[20, 29, 41] in other literature and has demonstrated advantages in the Prompt Engineering domain. However, as pointed out in [45], this type of method is ineffective in LLM-jailbreak because the obtained embeddings (i.e., soft prompts) often have no corresponding token. In our experiments, we reached a similar conclusion to [45], where the PEZ method completely failed to achieve any successful jailbreaks, i.e., $\text{ASR}\approx 0$ for all classes. Thus, the results of PEZ are omitted from Tables 1 and 2. The Soft Prompting (SoftP) method [29] achieves slightly better performance than PEZ; however, its ASR remains lower than that of token-based methods, as shown in Tables 1 and 2.
We argue that this failure stems from the absence of constraints during the embedding optimization process. Without such constraints, the optimized embeddings tend to deviate significantly from the distribution of natural text embeddings, making it difficult to be converted back to text tokens. In contrast, our approach directly optimizes images, which are processed through the visual module. This ensures that the resulting visual embeddings inherently remain closer to the distribution of text embeddings. As a result, when projecting these visual embeddings into text tokens, the projection error is significantly reduced.
To validate our analysis, we quantitatively measure the error incurred when converting optimized embeddings back to tokens. For both the Soft Prompting method and our approach, after converting JBemb $e^{k}$ to JBtxt via nearest neighbor search, we re-embed the obtained JBtxt to produce $\hat{e}^{k}$. We then calculate the cosine similarity between $\hat{e}^{k}$ and $e^{k}$. A higher similarity score indicates a lower conversion error. In our experiments, the average similarity for soft prompting is $\text{Sim}_{sp}=0.17$, whereas our method achieves a substantially higher average similarity of $\text{Sim}_{ours}=0.63$. These results compellingly corroborate our analysis.
4.1.3 Comparison on HarmBench
Besides the AdvBench-M dataset, we also evaluate jailbreaking performance on the HarmBench dataset [25]. From Table 3, our approach significantly outperforms all baseline methods. Compared to AdvAttack, which shows strong performance on HarmBench, our method achieves a 4.5% improvement on LLaMA-2 (72.0% vs. 67.5% ASR on the entire dataset) and a 1.8% improvement on InternLM-2.5 (63.0% vs. 61.2% ASR), fully demonstrating the effectiveness of our approach.
Table 4: Black-box Jailbreak of different LLMs via InternVL2.0-Jailbreak.
Class Black-box
Mistral-v0.3 Gemma2 Deepseek-R1 Grok-3 DouBao1.5
Unlawful violence 72.2 88.9 72.2 38.9 5.6
Financial crimes 96.4 92.8 89.2 62.5 39.2
Property crimes 93.7 92.1 78.1 67.2 31.2
Drug crimes 86.7 80.0 86.7 26.7 6.7
Weapons crimes 88.9 44.4 77.8 5.57 11.1
Cyber crimes 87.5 85.0 87.5 77.5 35.0
Hate 66.7 53.3 40.0 60.0 13.3
Suicide/Self-Harm 22.2 61.1 16.7 66.7 0.0
Fake info 80.0 93.3 86.7 33.3 40.0
Average ASR 77.1 76.7 70.5 48.7 20.1
4.2 Black-box Jailbreak
For the black-box jailbreaking scenario, we generate the JBtxt using a surrogate LLM and transfer it to jailbreak the target LLM. In practice, we first construct an MLLM according to the surrogate LLM and then perform jailbreaking on this surrogate MLLM to obtain the JBtxt.
In our experiments, we selected InternLM2.5 as the surrogate LLM. In practice, we employed InternVL2.0 as the corresponding MLLM, as it contains a InternLM2.5 inside. We then performed jailbreaking on InternVL2.0 to generate the JBtxt. Finally, this JBtxt was transferred to jailbreak target models, including Mistral-7B-v0.3, Gemma2-7B, Deepseek-R1, Grok-3, and DouBao1.5. The results are reported in Table 4. The experimental results demonstrate our approach has a strong black-box jailbreaking performance. We achieved remarkably high ASR of 77.1%, 76.7%, and 70.5% against Mistral-v0.3, Gemma2, and Deepseek-r1, respectively. However, we observed comparatively moderate efficacy when targeting proprietary models, with ASR of 48.7% on Grok-3 and 20.1% on DouBao1.5, suggesting these proprietary models incorporate more robust defensive mechanisms against jailbreaking.
Table 5: Image-text Semantic Matching. We compare three initialization schemes.
Initialization Clip-score MLLM LLM
$D_{train}$ $D_{test}$ $D_{test}^{others}$ $D_{train}$ $D_{test}$ $D_{test}^{others}$
Random-based 0.1473 73.32 72.03 73.42 80.74 74.88 75.62
Ranking-based 0.2083 75.18 76.75 75.44 82.22 77.07 75.94
Ours 0.5773 83.70 80.13 76.61 87.40 85.03 78.99
4.3 Discussion
4.3.1 Image-text Semantic Matching
In our approach, MLLM-jailbreak is achieved by perturbing the JBinit with a perturbation $\Delta$, i.e., $\text{JBimg}=\text{JBinit}+\Delta$. In this section, we illustrate that finding an appropriate JBinit is crucial for both successful MLLM-jailbreak and subsequent LLM-jailbreak.
In our experiments, we compare three distinct schemes for finding JBInit. The first scheme involves randomly sampling one image from all images in the AdvBench-M dataset, regardless of its harmful class. The second scheme restricts the sampling to images within the same harmful class as the query, and further employs CLIP ViT-L/14 to identify the image that best aligns with the semantics of the corresponding class. We refer to this as the ranking-based scheme. The third scheme is our proposed image-text semantic matching approach.
We compare the three initialization schemes in Table 5. In terms of CLIP similarity score, the ranking-based scheme demonstrates a marginal improvement over the random sampling scheme. In contrast, our image-text matching scheme significantly outperforms both baselines, yielding substantially higher CLIP scores. This improvement stems from the fact that our method explicitly and directly optimizes the JBInit to maximize alignment with harmful queries in the CLIP embedding space.
In terms of jailbreaking ASR, we observe that selecting an appropriate JBInit significantly boosts both MLLM-jailbreaking and LLM-jailbreaking ASR. Notably, as the CLIP similarity score increases, the ASR consistently improves. We argue that the underlying reason is that when the embedding of JBInit is closely aligned with that of the harmful queries, JBInit can significantly influence the LLM’s response generation—effectively turning refusal responses like “Sorry, I cannot” into affirmative replies such as “Sure, here is,” thereby enabling a successful jailbreak.
Figure 3: Cross-class generalization for LLM-jailbreak. For example, the first row illustrates that we generate JBtxt on the “Violence" and then apply it to jailbreak other harmful classes.
4.3.2 Cross-class Generalization
In real-world jailbreak scenarios, it is desirable for the JBtxt generated for a specific harmful class to generalize and effectively jailbreak other harmful classes.
To evaluate this cross-class generalization capability, we conduct an experiment where JBtxt is generated based on a specific harmful class and subsequently tested on all remaining classes. For instance, after generating a JBtxt using Class 1’s training set $D_{train}^{1}$, we evaluate its ASR on the testing sets of the other classes: $D_{test}^{2}$, $D_{test}^{3}$, …, $D_{test}^{9}$. This setup allows us to assess whether a JBtxt optimized for one class can successfully jailbreak across different harm classes.
The results of the cross-class generalization experiment are presented in Fig.3. We observe that certain classes, such as “Financial Crimes", “Property Crimes", and “Cyber Crimes" exhibit strong generalizability. Specifically, the JBtxt generated from any harmful class can effectively jailbreak these three classes. In contrast, classes like “Hate" and “Suicide and Self-Harm" prove challenging to generalize. Notably, even when the JBtxt is generated from these two classes themselves, the testing ASR remains low, as reflected by the diagonal entries in Fig.3.
Additionally, we observe that cross-class generalization exhibits distinct correlations between different classes. Certain pairs of classes demonstrate notably stronger correlations than others. For example, the JBtxt generated for the class “Unlawful Violence" can effectively jailbreak the class “Fake Info," while the JBtxt derived from “Weapons Crimes" shows strong efficacy in jailbreaking “Drug Crimes." This phenomenon implies that to improve the ASR for a given class, it may be advantageous to leverage harmful queries not only from that specific class but also from its correlated classes. This approach is particularly valuable when collecting sufficient queries for a specific harmful class is difficult. In future work, we aim to investigate the underlying factors driving these correlations among harmful classes.
5 Conclusion
This paper introduces an efficient LLM-jailbreaking method by constructing a multimodal large language model and executing MLLM-jailbreak. Unlike token-based jailbreaking techniques, our approach is notably efficient (approximately 3% of GCG’s runtime), as it exploits vulnerabilities within the visual module of the MLLM. Compared to embedding-based methods, we employ the visual module as a regularizer, ensuring that our jailbreaking embedding correspond to valid tokens. Furthermore, our approach exhibits superior cross-class generalization ability. Importantly, our indirect jailbreaking scheme provides notable flexibility for black-box jailbreak, enabling successful jailbreaks of recent LLMs such as Mistral-v0.3, Gemma2, Deepseek-R1, Grok-3, and DouBao1.5.